# Quantum variational calculus on a lattice

### v2

@@ 3.2. Euler-Lagrange equations
which we see is $\geq 0$ and zero only for the zero solution. The mix of sine and cosine modes depends on the parameters and here is instructive to take them to be the values of $q(0)$ and $E(q)$. One can analyse this analytically or just go back to the equations of motion, solved by recursion with
$q(1)={q(0)\over 2}\left(2-\omega^{2}+\sqrt{8{E(q)\over mq(0)^{2}}-4\omega^{2}+\omega^{4}}\right)$
@@
(19) $E[q]\geq E_{c}:={m\over 8}\omega^{2}(4-\omega^{2})q(0)^{2}.$
This is in line with classical mechanics where the total energy $E$ cannot be less than the potential energy at any point (here at $i=0$). For example, we can take $q(0)=0$ then $E_{c}=0$ and we see only sine solutions increasing in amplitude with increasing $E>0$ according to $E={1\over 8}m\omega^{2}(4-\omega^{2})q_{2}^{2}$ from (18). This modifies the relation $E={1\over 2}m\omega^{2}q_{2}^{2}$ in classical mechanics, parallel to the familiar fact in lattice field theory that the dispersion relation gets modified when $\omega$ is not small. Also parallel to the field theory case, if we let $\omega^{2}$ be negative then there are (real) unbounded exponential solutions. As we increase $\omega$, the amplitude for a given energy decreases until a minimum at $\omega=\sqrt{2}$, while the period $2\pi/\omega^{\prime}$ shrinks towards period 2, i.e alternating and blowing up in amplitude as $\omega\to 2$ from below. The case $\omega>2$ is not physical (if the discretisation is a model of Planck scale effects then these are transplankian modes) and here solutions blow up exponentially while alternating in $i$ (the ‘maximum frequency’). These effects are illustrated in Figure 1. Note that $E<E_{c}$ leads to complex oscillatory solutions but these are not part of our theory, and indeed the conserved energy for the complex case (which would be relevant to the scalar field theory point of view) is different as it involves conjugation of one of the fields.
@@
We then coordinatize the fibre of $J^{\infty}$ by $u_{I}$ corresponding to the coefficient of $e^{I}$ and have $C(J^{\infty})=C(G)[u_{I}]$ as before. The evaluation map is $e_{\infty}:G\times F\to J^{\infty}$ where $F$ is the space of fields (for a rank 1 real bundle it is just another copy of $C(G)$ up to completions) and pulling back gives $e_{\infty}^{*}:\Omega(J^{\infty})\to\Omega(G)\underline{\otimes}\Omega(F)$. Assuming (for the purposes of deriving the formulae) that this is an inclusion gives the same results as Theorem 3.1. The only difference is that $\mathbb{Z}$ is replaced by $G$ and now $a\in\mathcal{C}$ as discussed rather than $\pm$ as before.
Next, for the Euler-Lagrange equations, we still have a unique (up to scale) top form ${\rm Vol}=e^{1}\wedge\cdots\wedge e^{|\mathcal{C}|}$ and we use this to define $\mathrm{Vol}_{a}$ as a product of the 1-forms without $e^{a}$ such that $\mathrm{Vol}=e^{a}\wedge\mathrm{Vol}_{a}$ (no sum). Then the computation of $\mathrm{d}_{V}(L\mathrm{Vol})$ and hence of the EL form and boundary form goes as in Theorem 3.2, wielding the same results, now with $a\in\mathcal{C}$. Similarly, Section 3.3 can be reproduced up to the introduction of a Lagrangian and correspondingly the form of the stress-energy tensor $T_{ab}$, but the computation of the conserved charges needs extra care, and is also a little different when we make the interpretation 1+0 classical field theory. We will give the results in detail for $\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$, but the method works in the same way for any Abelian group of the form $\mathbb{Z}\times G$ or $\mathbb{Z}_{N}\times G$, replacing $\mathcal{C}_{m-1}$ as defined below by $\mathcal{C}_{G}$.
@@ 4.1. EL equations and translation symmetries for $X=\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$
We focus now on scalar field theory on $X=\mathbb{Z}^{m}$ with the Euclidean metric, meaning that $(e^{a},e^{b})=\delta^{a,b^{-1}}$. For the calculus, we take $\mathcal{C}$ to be the set of all the positive and negative directions in $\mathbb{Z}^{m}$
@@
where we again introduce the factor $1\over 2$} in front of the kinetic term in order to recover the classical action from Example 2.2. Recall that the minus cancels the one coming from $(\mathrm{d}\phi,\mathrm{d}\phi)$, and the $1\over 2$} is there since every classical direction corresponds to two directions in $\mathcal{C}$ (namely $e^{i\pm}$). Given the Lagrangian $L=\frac{1}{4}\sum_{a\in\mathcal{C}}u_{a}^{2}-\frac{1}{2}m^{2}u^{2}$ we can use Theorem 3.2 to compute the EL and boundary forms
@@
$D_{t}(j^{t}-R_{t^{-1}}j^{t^{-1}})+\sum_{x\in\mathbb{Z}_{m-1}}D_{x}(j^{x}-R_{x^{-1}}j^{x^{-1}}).$
For translation symmetries, since $j^{a}=\sum_{b\in\mathcal{C}}\epsilon^{b}T_{a^{-1}b}$ we have $|\mathcal{C}|/2$ conserved quantities due to 4.2, with the energy corresponding to the $t$ direction and momenta corresponding to the $\mathbb{Z}^{m-1}$ spatial directions being
@@
where for the momentum $P_{b}$ we had to use integration by parts and $\int_{\mathbb{Z}^{m-1}}\partial_{b}=0$ to arrive to $P_{b}$.
@@ 4.2. Scalar field theory in the (1+1)-dimensional case
We focus now on (1+1)-dimensional case of Euclidean field theory, where the Klein-Gordon equation $(\Delta+m^{2})\phi(t,x)=0$ is solved by the plane wave solutions
@@
satisfying as well known dispersion relation for a lattice
@@
as in [23, 25, 26]. This is normally justified from its appearance in poles in the propagator after Fourier transform to $\omega,\kappa$ as the Fourier conjugate variables, whereas our considerations here will be in line with the continuum version from classical variational calculus in Example 2.6, for which we have the required limit $m^{2}\simeq\omega^{2}+\kappa^{2}$ for $\omega,\kappa\ll 1$. Note that in Section 3.3, we saw that a classical particle $q$ in a quadratic potential $V(q)=\frac{1}{2}m\omega^{2}q^{2}$ only allowed for plane-wave solutions for $\omega<2$, with a different type of numerical solutions appearing for $\omega>2$. Here, we similarly find that the mass $m$ of the field for a plane wave (now playing the role of the frequency $\omega$ in the previous example) is also bounded as $m\leq 2\sqrt{2}$ as $\cos(\omega),\cos(\kappa)\in[-1,1]$.
@@
$m\mapsto m^{\prime}=\sqrt{8-m^{2}},\quad\omega\mapsto\omega^{\prime}=\pi-\omega,\quad\kappa\mapsto\kappa^{\prime}=\pi-\kappa$
@@
Figure 2. Plots of the dispersion relation, energy and momentum densities at masses $m=0.9,1.7,2,2.25,2.67$ and comparison with the continuum counterparts in the Euclidean case dashed. The $m>2$ values are the $m^{\prime}=\sqrt{8-m^{2}}$ counterparts of the $m<2$ values but with 0.01-0.01 offset to separate curves in the last plot. For small masses, the lattice and continuum plots almost match. At $m=2$, waves with all $\kappa\in[-\pi,\pi]$ propagate in the lattice. As we approach $m=2\sqrt{2}$, waves closer to $\kappa=0$ stop propagating first, until eventually at $m=2\sqrt{2}$, only waves with $\kappa=-\pi,\pi$ propagate. The lower right plot shows the energy density against the momentum density for $A=\sqrt{2/\langle\mathcal{E}_{dens}[\phi]\rangle}$ as in Equation (24).
@@
$=\frac{A^{2}}{8}\left(4\cos^{2}(\kappa x+t\omega)-\cos(2\kappa(x-1)+2t\omega)-\cos(2(\kappa x+\kappa+t\omega))-2\cos(2\omega)\right)$
$\mathcal{P}_{x,dens}[\phi]$ $={1\over 2}(\partial_{t^{-1}}\phi)\left(\partial_{x^{-1}}\phi-\partial_{x}\phi\right)=A^{2}\sin(\kappa)\sin(\omega)\sin^{2}\left(\kappa x+\frac{\kappa}{2}+t\omega-\frac{\omega}{2}\right)$
@@
(22) $\langle\mathcal{E}_{dens}[\phi]\rangle\in\left[0,\frac{A^{2}}{2}\right]\quad\quad\langle\mathcal{P}_{x,dens}[\phi]\rangle\in\left[-\frac{A^{2}}{2},\frac{A^{2}}{2}\right]$
due to lattice effects. These densities and their comparison with the classical expressions are also shown in Figure 2. The symmetry between $m$ and $m^{\prime}=\sqrt{8-m^{2}}$ is also present in the energy and momentum density, where the plots for $m=0.9,1.7$ centred around $\kappa$ are mirrored by plots for $m^{\prime}=2.25,2.67$ now centered around $\kappa=\pi$. We did not take exactly $\sqrt{8-m^{2}}$ so that the curves would not be exactly on top of each other in the energy-momentum plot.
@@
This relation results in the energy against momentum plot in Figure 2, where the $s_{\kappa}=+1$ (small $\kappa\ll 1$) solution is comparable to the classical case $\langle\mathcal{E}_{x,dens}[\phi]\rangle=\sqrt{m^{2}-\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}}$ shown dashed. The solutions (24) again display the $m,\,m^{\prime}=\sqrt{8-m^{2}}$ symmetry in that $\langle\mathcal{E}_{dens}[\phi]\rangle_{\pm}$ for $m$ correspond to $\langle\mathcal{E}_{dens}[\phi]\rangle_{\mp}$ for $m^{\prime}$.
In this case, we want to construct a stress-energy tensor $T$ which satisfies the divergence free condition for the Minkoswki metric, where
$j=\sum_{a,b\in\mathcal{C}}\epsilon^{b}T^{a}{}_{b}\mathrm{Vol}_{a}$
$T_{ab}=\sum_{c\in\mathcal{C}}\eta_{ac}T^{c}{}_{b}=\eta^{a}T^{a^{-1}}{}_{b}$.
$T_{ab}=-\left(\frac{1}{2}u_{a}u_{b}+\frac{1}{4}u_{ab}u_{b}+\frac{\delta^{a^{-1}}_{b}}{\eta^{a}}L\right)$
Figure 4. Dipersion relation, energy density and momentum densities in the Minkowski case for the masses $m=0.5,\sqrt{2},1.9$. At lower masses, $\kappa=0$ is a local minimum of the energy density, while it is a local maximum for waves with higher masses. In the lower right, the energy density is plotted against the momentum density following Equation (28).
@@
as also found in [14]. This recovers the classical limit $m^{2}=\omega^{2}-\kappa^{2}$ when $\omega,\kappa\ll 1$, i.e. much smaller than the lattice spacing. We find that in this case the mass is bounded as $m<2$ for physical states, and find that the frequency behaves as
$\omega(\kappa)=\arccos\left(\cos(\kappa)-{m^{2}\over 2}\right)$
@@
$\mathcal{P}_{x,dens}[\phi]$ $=-A^{2}\sin(\kappa)\sin(\omega)\sin^{2}\left(\left(t-\frac{1}{2}\right)\omega-\frac{1}{2}\kappa(2x+1)\right)$
@@
as before, here with a minus in the momentum density due to the metric signature. As expected, these are time independent and tend to the continuum values for small $\omega,\kappa$. They are also plotted in Figure 4. We see that the profile of the energy density changes depending on whether $m<\sqrt{2}$, $m=\sqrt{2}$ or $m>\sqrt{2}$. In the first case, there are two local maxima and 3 local minima, one at $\kappa=0$ and two at the edges of the domain. For $m>\sqrt{2}$, the point at $\kappa=0$ is a local maximum. Comparing these with the plot of the momentum density, we find that these energy minima, either at $\kappa=0$ or at the edges of the $\kappa$-domain, always correspond to plane waves with zero momentum density.
Figure 5. Phase and group velocities in the Minkwoski case for $m=0.5,\sqrt{2},1.9$ and comparison with their classical counterparts shown dashed.
@@
$\approx\left(1+\frac{m^{2}}{2}\right)\langle\mathcal{E}_{dens}\rangle^{2}-\left(1+{m^{2}\over 4}\right)\frac{A^{2}}{2}\langle\mathcal{E}_{dens}\rangle\,m^{2}$
@@
as illustrated in Figure 4 against their classical counterparts. For small $\kappa$, the relevant solution that deforms the classical case is with $s_{\kappa}=+1$ as in the introduction.

### v3

@@ 3.2. Euler-Lagrange equations
which we see is $\geq 0$, and zero only for the zero solution. The mix of sine and cosine modes depends on the parameters and here is instructive to take them to be the values of $q(0)$ and $E[q]$. One can analyse this analytically or just go back to the equations of motion, solved by recursion with
$q(1)={q(0)\over 2}\left(2-\omega^{2}+\sqrt{8{E[q]\over mq(0)^{2}}-4\omega^{2}+\omega^{4}}\right)$
@@
(17) $E[q]\geq E_{c}:={m\over 8}\omega^{2}(4-\omega^{2})q(0)^{2}=\left(1-\frac{\omega^{2}}{4}\right)V(q(0)).$
In classical mechanics, the total energy cannot be less than the potential energy and here, we find a deformation of this fact (here at $i=0$) as recovered for small $\omega$. For example, we can take $q(0)=0$ then $E_{c}=0$ and we see only sine solutions increasing in amplitude with increasing $E>0$ according to $E={1\over 8}m\omega^{2}(4-\omega^{2})q_{2}^{2}$ from (16). This modifies the relation $E={1\over 2}m\omega^{2}q_{2}^{2}$ in classical mechanics, parallel to the familiar fact in lattice field theory that the dispersion relation gets modified when $\omega$ is not small. Also parallel to the field theory case, if we let $\omega^{2}$ be negative then there are (real) unbounded exponential solutions. As we increase $\omega$, the amplitude for a given energy decreases until a minimum at $\omega=\sqrt{2}$, while the period $2\pi/\omega^{\prime}$ shrinks towards period 2, i.e alternating and blowing up in amplitude as $\omega\to 2$ from below. The case $\omega>2$ is not physical (if the discretisation is a model of Planck scale effects then these are transplankian modes) and here solutions blow up exponentially while alternating in $i$ (the ‘maximum frequency’). These effects are illustrated in Figure 1. Note that $E<E_{c}$ leads to complex oscillatory solutions but these are not part of our theory, and indeed the conserved energy for the complex case (which would be relevant to the scalar field theory point of view) is different as it involves conjugation of one of the fields.
Figure 1. Discrete time harmonic oscillator for $m=1$ and energy $E=1$ with increasing $\omega$. For the sine modes with $\omega<2$, the dashed line is the amplitude $q_{2}$ such that $E={1\over 8}m\omega^{2}(4-\omega^{2})q_{2}^{2}$, with minimum amplitude for $\omega=\sqrt{2}$. The period $2\pi/\omega^{\prime}$ is significantly smaller than $2\pi/\omega$ as $\omega$ increases. As $\omega\to 2$ from below, the amplitude blows up and the period tends towards the minimum alternating case. We also see a beat frequency emerging. The $\omega>2$ case has alternating exponential growth. Plots have been smoothly interpolated for visualization.
@@
We then coordinatise the fibre of $J^{\infty}$ by $u_{I}$ corresponding to the coefficient of $e^{I}$ and have $C(J^{\infty})=C(G)[u_{I}]$ as before. The evaluation map is $e_{\infty}:G\times F\to J^{\infty}$ where $F$ is the space of fields (for a rank 1 real bundle it is just another copy of $C(G)$ up to completions) and pulling back gives $e_{\infty}^{*}:\Omega(J^{\infty})\to\Omega(G)\underline{\otimes}\Omega(F)$. Assuming (for the purposes of deriving the formulae) that this is an inclusion gives the same results as Theorem 3.1. The only difference is that $\mathbb{Z}$ is replaced by $G$ and now $a\in\mathcal{C}$ as discussed rather than $\pm$ as before.
Next, for the Euler-Lagrange equations, we still have a unique (up to scale) top form ${\rm Vol}=e^{1}\wedge\cdots\wedge e^{|\mathcal{C}|}$ and we use this to define $\mathrm{Vol}_{a}$ as a product of the 1-forms without $e^{a}$ such that $\mathrm{Vol}=e^{a}\wedge\mathrm{Vol}_{a}$ (no sum). Then the computation of $\mathrm{d}_{V}(L\mathrm{Vol})$ and hence of the EL form and boundary form goes as in Theorem 3.2, yielding the same results, now with $a\in\mathcal{C}$. Similarly, Section 3.3 can be reproduced up to the introduction of a Lagrangian and correspondingly the form of the stress-energy tensor $T_{ab}$, but the computation of the conserved charges needs extra care, and is also a little different when we make the interpretation 1+0 classical field theory. We will give the results in detail for $\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$, but the method works in the same way for any Abelian group of the form $\mathbb{Z}\times G$ or $\mathbb{Z}_{N}\times G$, replacing $\mathcal{C}_{m-1}$ as defined below by $\mathcal{C}_{G}$.
@@ 4.1. EL equations and translation symmetries for $X=\mathbb{Z}^{m},\mathbb{Z}^{1,m-1}$
We now focus on scalar field theory on $X=\mathbb{Z}^{m}$ with the Euclidean metric, meaning that $(e^{a},e^{b})=\delta^{a,b^{-1}}$. For the calculus, we take $\mathcal{C}$ to be the set of all the positive and negative directions in $\mathbb{Z}^{m}$
@@
where we again introduce the factor $1\over 2$} in front of the kinetic term in order to recover the classical action from Example 2.2. Recall that the minus cancels the one coming from $(\mathrm{d}\phi,\mathrm{d}\phi)$, and the $1\over 2$} is there since every classical direction corresponds to two directions in $\mathcal{C}$ (namely $e^{i\pm}$). Given the Lagrangian $L=\frac{1}{4}\sum_{a\in\mathcal{C}}u_{a}^{2}-\frac{1}{2}m^{2}u^{2}$, and we can use Theorem 3.2 to compute the EL and boundary forms
@@
$D_{t}(j^{t}-R_{t^{-1}}j^{t^{-1}})+\sum_{x}D_{x}(j^{x}-R_{x^{-1}}j^{x^{-1}}).$
where $x$ are the spatial directions in $\mathbb{Z}^{m-1}$. ∎
For translation symmetries, we have $|\mathcal{C}|/2$ conserved quantities due to 4.2 as $j^{a}=\sum_{b\in\mathcal{C}}\epsilon^{b}T_{a^{-1}b}$, with the energy corresponding to the $t$ direction and momenta corresponding to the $\mathbb{Z}^{m-1}$ spatial directions being
@@
where for the momentum $P_{b}$ we had to use integration by parts and $\int_{\mathbb{Z}^{m-1}}(\partial_{b}f)\mathrm{Vol}_{Z^{m-1}}=0$ to arrive to $P_{b}$.
@@
In this case, we want to construct a stress-energy tensor $T$ which satisfies the divergence free condition for the Minkowski metric, where
$j=\sum_{a,b\in\mathcal{C}}\epsilon^{b}T^{a}{}_{b}\mathrm{Vol}_{a}$
$T_{ab}=\sum_{c\in\mathcal{C}}\eta_{ac}T^{c}{}_{b}=\eta^{a}T^{a^{-1}}{}_{b}$,
$T_{ab}=-\left(\frac{1}{2}u_{a}u_{b}+\frac{1}{4}u_{ab}u_{b}+\frac{\delta^{a^{-1}}_{b}}{\eta^{a}}L\right).$
@@ 4.2. Scalar field theory in the (1+1)-dimensional case
We focus now on (1+1)-dimensional case of scalar field theory, starting with the Euclidean version where the Klein-Gordon equation $(\Delta+m^{2})\phi(t,x)=0$ is solved by the plane wave solutions
@@
satisfying the following dispersion relation for a lattice
@@
as in [23, 25, 26]. This is normally justified from its appearance in poles in the propagator after Fourier transform to $\omega,\kappa$ as the Fourier conjugate variables, whereas our considerations here will be in line with the continuum version from classical variational calculus in Example 2.6, for which we have the required limit $m^{2}\simeq\omega^{2}+\kappa^{2}$ for $\omega,\kappa\ll 1$ much smaller than the lattice spacing. Note that in Section 3.3, we saw that a classical particle $q$ in a quadratic potential $V(q)=\frac{1}{2}m\omega^{2}q^{2}$ only allowed for plane-wave solutions for $\omega<2$, with a different type of numerical solutions appearing for $\omega>2$. Here, we similarly find that the mass $m$ of the field for a plane wave (now playing the role of the frequency $\omega$ in the previous example) is also bounded as $m\leq 2\sqrt{2}$ as $\cos(\omega),\cos(\kappa)\in[-1,1]$.
@@
$m\mapsto m^{\prime}=\sqrt{8-m^{2}},\quad\omega\mapsto\omega^{\prime}=\pi-\omega,\quad\kappa\mapsto\kappa^{\prime}=\pi-\kappa,$
@@
Figure 2. Plots of the dispersion relation, energy and momentum densities at masses $m=0.9,1.7,2,2.25,2.67$ and comparison with the continuum counterparts in the Euclidean case, shown as dashed. The $m>2$ values are the $m^{\prime}=\sqrt{8-m^{2}}$ counterparts of the $m<2$ values but with 0.01-0.01 offset to separate curves in the last plot. For small masses, the lattice and continuum plots almost match. At $m=2$, waves with all $\kappa\in[-\pi,\pi]$ propagate in the lattice. As we approach $m=2\sqrt{2}$, waves closer to $\kappa=0$ stop propagating first, until eventually at $m=2\sqrt{2}$, only waves with $\kappa=-\pi,\pi$ propagate. The lower right plot shows the energy density against the momentum density for $A=\sqrt{2/\langle\mathcal{E}_{dens}[\phi]\rangle}$ as in equation (23).
@@
$=\frac{A^{2}}{8}\left(4\cos^{2}(\kappa x+t\omega)-\cos(2\kappa(x-1)+2t\omega)-\cos(2(\kappa x+\kappa+t\omega))-2\cos(2\omega)\right),$
$\mathcal{P}_{x,dens}[\phi]$ $={1\over 2}(\partial_{t^{-1}}\phi)\left(\partial_{x^{-1}}\phi-\partial_{x}\phi\right)=A^{2}\sin(\kappa)\sin(\omega)\sin^{2}\left(\kappa x+\frac{\kappa}{2}+t\omega-\frac{\omega}{2}\right).$
@@
(21) $\langle\mathcal{E}_{dens}[\phi]\rangle\in\left[0,\frac{A^{2}}{2}\right],\quad\quad\langle\mathcal{P}_{x,dens}[\phi]\rangle\in\left[-\frac{A^{2}}{2},\frac{A^{2}}{2}\right],$
due to lattice effects. These densities and their comparison with the classical expressions are also shown in Figure 2, where we have scaled these quantities by $1/A^{2}$. The symmetry between $m$ and $m^{\prime}=\sqrt{8-m^{2}}$ is also present in the energy and momentum density, where the plots for $m=0.9,1.7$ centred around $\kappa$ are mirrored by plots for $m^{\prime}=2.25,2.67$ now centered around $\kappa=\pi$. We did not take exactly $\sqrt{8-m^{2}}$ so that the curves would not be exactly on top of each other in the energy-momentum plot.
@@
We focus on the positive energy solutions but there is a mirror image of negative ones for the other choice of sign of the outer square root. This relation results in the energy against momentum plot in Figure 2, where the $s_{\kappa}=+1$ (small $\kappa\ll 1$) solution is comparable to the classical case $\langle\mathcal{E}_{x,dens}[\phi]\rangle=\sqrt{m^{2}-\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}}$ shown dashed. The solutions (23) again display the $m,\,m^{\prime}=\sqrt{8-m^{2}}$ symmetry in that $\langle\mathcal{E}_{dens}[\phi]\rangle_{\pm}$ for $m$ correspond to $\langle\mathcal{E}_{dens}[\phi]\rangle_{\mp}$ for $m^{\prime}$.
Figure 4. Dispersion relation, energy density and momentum densities in the Minkowski case for the masses $m=0.5,\sqrt{2},1.9$. At lower masses, $\kappa=0$ is a local minimum of the energy density, while it is a local maximum for waves with higher masses. In the lower right, the energy density is plotted against the momentum density following equation (26).
@@
as also found in [14]. This recovers the classical limit $m^{2}=\omega^{2}-\kappa^{2}$ when $\omega,\kappa\ll 1$, i.e. much smaller than the lattice spacing. We find that in this case the mass is bounded as $m<2$ for physical states, and that the frequency behaves as
$\omega(\kappa)=\arccos\left(\cos(\kappa)-{m^{2}\over 2}\right),$
@@
$\mathcal{P}_{x,dens}[\phi]$ $=-A^{2}\sin(\kappa)\sin(\omega)\sin^{2}\left(\left(t-\frac{1}{2}\right)\omega-\frac{1}{2}\kappa(2x+1)\right),$
@@
as before, here with a minus in the momentum density due to the metric signature. As expected, these are time independent and tend to the continuum values for small $\omega,\kappa$. They are also plotted in Figure 4 over the square amplitude $A^{2}$. We see that the profile of the energy density changes depending on whether $m<\sqrt{2}$, $m=\sqrt{2}$ or $m>\sqrt{2}$. In the first case, there are two local maxima and 3 local minima, one at $\kappa=0$ and two at the edges of the domain. For $m>\sqrt{2}$, the point at $\kappa=0$ is a local maximum. Comparing these with the plot of the momentum density, we find that these energy minima, either at $\kappa=0$ or at the edges of the $\kappa$-domain, always correspond to plane waves with zero momentum density.
Figure 5. Phase and group velocities in the Minkowski case for $m=0.5,\sqrt{2},1.9$ and comparison with their classical counterparts shown dashed.
@@
$\approx\left(1+\frac{m^{2}}{2}\right)\langle\mathcal{E}_{dens}\rangle^{2}-\left(1+{m^{2}\over 4}\right)\frac{A^{2}}{2}\langle\mathcal{E}_{dens}\rangle\,m^{2},$
@@
as illustrated in Figure 4 against their classical counterparts. We again focus on the positive energy solutions but there is a mirror image of negative energy ones. For small $\kappa$, the relevant solution that deforms the classical case is with $s_{\kappa}=1$ as in the introduction. In this case, we find that the configuration with vanishing momentum density $\langle\mathcal{P}_{x,dens}[\phi]\rangle=0$ corresponds to a local minima in the energy for $m<\sqrt{2}$ which is global for $m\leq 1$, and a global maximum for $m\geq\sqrt{2}$. Note that the plots also show the $s_{\kappa}=-1$ solutions which are typically of lower energy and connect through zero to the negative energy solutions. These modes do not have a continuum counterpart and it is unclear as to their physical role. It is interesting that even the $s_{\kappa}=+1$ solutions have lowest energy at nonzero spatial momentum when $1<m<\sqrt{2}$, namely at $|\langle\mathcal{P}_{x,dens}[\phi]\rangle|=1$.